OpenStudy (anonymous):

What is the molecular mass (in g/mol) of 5.4 g of a gas at 1.8 atm and 4.3 oC if the gas occupies a 2.1 L vessel?

OpenStudy (gebooors):

Use equation pV = nRT and solve amount of substance n. Then you can solve molar mass M using n= m/M Which you can convert M = m/n.